The stock of a Thai based company is currently trading at 50,000 Bhat. Your financial advisor tells you that she expects the stock to pay a dividend of 5,000 Bhat per share, and her best forecast for that same stock, 1 year from now is 47,000 Bhat

You are in Malaysia with bank deposit that pays 8% rate of interest. If the Bhat is currently trading at 0.12 Ringgit, and you expect it will trade a 0.13 Ringgit 1 year from now, should you take some of your money out of the Malaysian bank and invest it with this company?
Answer #1

Baht is trading at 0.12 Ringgit. That means 1 baht is equal to 0.12 Ringgit, then 50000 baht will be = 6000 Ringgit. If that money is invested in stock and if one sells the stock after one year and earns dividend on it than. 47,000+5000 = 52,000. If one converts that into ringgit after one year than 1 baht = 0.13 Ringgit, 52,000 Baht = 52,000×0.13 = 6760.

So the 6000 which I invested, in the stock if I had invested in the local bank than I would have earned, 8% interest which goes to 6480 ringgits ( 6000*1.08)

So one should take money out of the bank as the return is more by investing in the stock however marginal the return per stock. One earns 6760-6480= 280 per stock.